Solve. dx/dt=-2x+7y,dy/dt=3x+2ysubject to the condition x(0)=9 and y(0)=-1
1
Expert's answer
2020-12-21T19:42:02-0500

Let us solve the sysytem:


"\begin{cases} \frac{dx}{dt}=-2x+7y \\ 
\frac{dy}{dt}=3x+2y \end{cases}" "x(0)=9" and "y(0)=-1"


The second eqution is equivalent to "x=\frac{1}{3}( \frac{dy}{dt}-2y)", and therefore, "\frac{dx}{dt}=\frac{1}{3}( \frac{d^2y}{dt^2}-2\frac{dy}{dt})" . Let us put these in the first equation:


"\frac{1}{3}( \frac{d^2y}{dt^2}-2 \frac{dy}{dt})=-2\cdot\frac{1}{3}( \frac{dy}{dt}-2y)+7y"


"\frac{1}{3} \frac{d^2y}{dt^2}-\frac{2}{3} \frac{dy}{dt}=-\frac{2}{3}\frac{dy}{dt}+\frac{4}{3}y+7y"


"\frac{1}{3} \frac{d^2y}{dt^2}=\frac{25}{3}y"


"\frac{d^2y}{dt^2}-25y=0"


The characteristic equation "k^2-25=0" of the last differential eqution has the solutions "k=5" and "k=-5." Consequently, its general solution is of the form


"y(t)=C_1e^{5t}+C_2e^{-5t}"


"\frac{dy}{dt}=5C_1e^{5t}-5C_2e^{-5t}"


Thus, "x=\frac{1}{3}( \frac{dy}{dt}-2y)=\frac{1}{3}( 5C_1e^{5t}-5C_2e^{-5t}-2(C_1e^{5t}+C_2e^{-5t}))=C_1e^{5t}-\frac{7}{3}C_2e^{-5t}."


If "x(0)=9" and "y(0)=-1", then "9=x(0)=C_1-\frac{7}{3}C_2" and "-1= y(0)=C_1+C_2" .


It follows that "C_1=2,\ C_2=-3."


Therefore, the solution is of the form:


"\begin{cases} x(t)=2e^{5t}+7e^{-5t}\\
y(t)=2e^{5t}-3e^{-5t}
\end{cases}"